# Pressure and strain effects on the ab initio electronic structure of La3Ni2O7

### v1

@@ Appendix H Bloch state plots
The lower and upper Ni-3 bands are sometimes respectively labeled as bonding and anti-bonding, according to a chemical convention. For these two states we plot in Fig.Β 17 the iso-surfaces of the Bloch wave functions at the X -point which lies in the corner of the BZ. As we can see, these names are not appropriated, the upper state being more βbondingβ than the lower one, with more weight on orbitals of inner oxygen atoms lying in between NiO2 planes.
Another Bloch state worth to be shown is the one corresponding to the La-5 band, which is the most sensitive to pressure, and in particular at the M point where it reaches its minimum. In Fig.Β 18 we see that these electrons are delocalized in the spacer, which explain the highly dispersive shape of the corresponding band. From 0 to 29.5Β GPa, this state is not occupied. It starts to be occupied beyond 29.5Β GPa, inducing self-doping of the NiO2 planes.

### v2

@@ Appendix I Bloch state plots
The lower and upper Ni-3 bands are sometimes respectively labeled as bonding and antibonding, according to a chemical convention. For these two states we plot in Fig. 18 the iso-surfaces of the Bloch wave functions at the X -point which lies in the corner of the BZ. As we can see, these names are not appropriated, the upper state being more “bonding” than the lower one, with more weight on orbitals of inner oxygen atoms lying in between NiO2 planes.
Another Bloch state worth to be shown is the one corresponding to the La-5 band, which is the most sensitive to pressure, and in particular at the M point where it reaches its minimum. In Fig. 19 we see that these electrons are delocalized in the spacer, which explain the highly dispersive shape of the corresponding band. From 0 to 29.5 GPa, this state is not occupied. It starts to be occupied beyond 29.5 GPa, inducing self-doping of the NiO2 planes.